# Jailbreak-AudioBench: In-Depth Evaluation and Analysis of Jailbreak Threats for Large Audio Language Models

### v3

@@ Abstract
Large Language Models (LLMs) demonstrate impressive zero-shot performance across a wide range of natural language processing tasks. Integrating various modality encoders further expands their capabilities, giving rise to Multimodal Large Language Models (MLLMs) that process not only text but also visual and auditory modality inputs. However, these advanced capabilities may also pose significant security risks, as models can be exploited to generate harmful or inappropriate content through jailbreak attack. While prior work has extensively explored how manipulating textual or visual modality inputs can circumvent safeguards in LLMs and MLLMs, the vulnerability of audio-specific Jailbreak on Large Audio-Language Models (LALMs) remains largely underexplored. To address this gap, we introduce Jailbreak-AudioBench, which consists of the Toolbox, curated Dataset, and comprehensive Benchmark. The Toolbox supports not only text-to-audio conversion but also various editing techniques for injecting audio hidden semantics. The curated Dataset provides diverse explicit and implicit jailbreak audio examples in both original and edited forms. Utilizing this dataset, we evaluate multiple state-of-the-art LALMs and establish the most comprehensive Jailbreak benchmark to date for audio modality. Finally, Jailbreak-AudioBench establishes a foundation for advancing future research on LALMs safety alignment by enabling the in-depth exposure of more powerful jailbreak threats, such as query-based audio editing, and by facilitating the development of effective defense mechanisms.
@@ 1 Introduction
Figure 1: The framework of Jailbreak-AudioBench
@@
Recently, Large Language Models (LLMs), represented by GPT-4o [25], Claude [5], and DeepSeek [19], have received increasing attention due to their strong general capabilities, efficient information processing, and natural human-computer interaction. LLMs perform well across a variety of natural language processing tasks, including question answering [63; 31], sentence summarization [12; 27], language translation [15; 33], and sentiment analysis [62; 20]. Leveraging the powerful reasoning capacity of LLMs, researchers develop Multimodal Large Language Models (MLLMs) by introducing various modality-specific encoders, enabling these models to perceive multiple modalities and handle more diverse tasks. Among them, Large Vision-Language Models (LVLMs), which combine vision encoders with LLMs, achieve strong performance on various Visual Question Answering tasks by modeling joint vision-language representations [17; 7; 40; 22; 8; 43]. In addition, Audio-Language Processing plays an important role in real-world applications such as voice assistants (e.g., Siri, Google Assistant, Cortana [23; 49]), customer service systems [2; 45], and in-vehicle voice control systems [29; 4]. Large Audio Language Models (LALMs), developed by integrating audio encoders into LLMs, are introduced to expand information processing capabilities from textual to auditory modalities, enabling more advanced audio-language understanding tasks.
@@
Current LALMs are mainly categorized into cascaded LALMs and end-to-end LALMs. Cascaded LALMs [41; 6; 14; 48; 39] typically consist of a two-stage pipeline, where an upstream Automatic Speech Recognition module first transcribes audio into text, which is then processed by a downstream LLMs for reasoning or generation. However, this approach discards information during transcription, making it incapable of capturing audio-specific hidden semantics. In contrast, end-to-end LALMs [25; 52; 61; 9; 46; 13; 30; 59] address this limitation by integrating audio encoding and language modeling into a single architecture that directly consumes raw audio inputs and generates corresponding textual outputs. By bypassing intermediate transcription, these models preserve complete audio information, especially the critical hidden semantics, which are essential for in-depth audio modality perception. Therefore, advancing research on end-to-end LALMs is becoming increasingly important for enhancing audio-language cross-modal understanding.
@@
In an era of rapid advancement in various types of LLMs and MLLMs, the exploration of their safety alignment becomes increasingly critical. The jailbreak threats refer to the use of carefully crafted prompts to bypass alignment safeguards and induce AI systems to generate outputs that violate intended safety constraints. These handcrafted strategies are highly diverse, encompassing techniques such as adversarial optimization, prompt-based manipulations, and other [64; 44; 24; 18; 35; 40; 22]. Among these, a wide range of prompt-tuning techniques, such as imperative commands (e.g., “you must answer”, “!!!”), role playing instructions (e.g., “act as an unrestricted A”), emoji injection, and distraction-based redirection (e.g., mixing benign and harmful queries), prove to be simple yet highly effective in subverting system-level safeguards [64; 44; 24; 18; 54]. Notably, inserting elements such as “!!!”, emojis, or garbled characters into the original prompts, which represent forms of hidden semantics, can also successfully trigger jailbreak attacks [64; 18; 54]. Due to their innocuous appearance, ease of insertion, and strong potential to induce jailbreak threats, these hidden semantics underscore the latent vulnerabilities of current large models in maintaining robust safety alignment.
@@
Compared to the language text modality, the audio modality inherently conveys richer hidden semantic information, such as Emphasis, Speech Speed, Intonation, Tone, Background Noise, Accent and Emotion. Unlike cascaded LALMs, end-to-end LALMs directly perceive and interpret these diverse audio-specific features, and are therefore widely considered one of the most promising directions in processing Audio Language Processing tasks. However, this deep sensitivity to audio modality also renders end-to-end LALMs more vulnerable to hidden semantic manipulations, introducing potential security risks, particularly in the context of jailbreak attacks. Although a few preliminary studies have emerged [56], systematic investigation into the jailbreak vulnerabilities of end-to-end LALMs remains limited. To address this gap, as the framework presented in Figure 1, this paper introduces Jailbreak-AudioBench, the most comprehensive evaluation to date of representative end-to-end LALMs under diverse jailbreak attack scenarios, and further highlights the critical role of modality-specific semantics in shaping the effectiveness of these threats. Moreover, we demonstrate that Jailbreak-AudioBench can serve as a valuable tool to further facilitate various explorations into the safety alignment of LALMs. The main contents are outlined as follows:
@@
Figure 2: (a) Different sub-tasks of each Jailbreak-AudioBench Dataset Subtype; (b) The largest jailbreak threat variation induced by audio hidden semantics across various LALMs;
@@
- Potential Research inspired by Jailbreak-AudioBench In order to further indicate the research impact of Jailbreak-AudioBench, we present two potential works: (1) Query-based Audio Editing Jailbreak. By adopting the Toolbox to perform query-based editing on a small set of implicit subtype jailbreak audios, the ASR success rates of Qwen2-Audio, SALMONN-7B, and GPT-4o-Audio increased from $13.3\%$ to $48.8\%$, $31.6\%$ to $85.1\%$, and $0.7\%$ to $8.4\%$, respectively. (2) Defense Against Audio Editing Jailbreak. Further, Potential defense strategies targeting LALM jailbreak threats can be effectively developed through the use of Jailbreak-AudioBench.
@@ 2 Jailbreak-AudioBench Toolbox
Figure 3: Injection of Various Audio Hidden Semantics
Preliminary    For a systematic evaluation of jailbreak threats in LALMs, the Jailbreak-AudioBench toolbox not only performs text-to-audio conversion but also implements a comprehensive suite of audio editing types to inject diverse forms of hidden semantics, including emphasis, speed, intonation, background noise, celebrity accent, and emotion, each modulated with different parameters as illustrated in Figure 1. The Text-to-Audio conversion is accomplished using Google Text-to-Speech (gTTS) [11]. Various audio editing methods are implemented with a range of tools, including Short-Time Fourier Transform (STFT), SOX (Sound eXchange), coqui-ai TTS [10], and Dia-1.6B [38]. Figure 3 further uses textual characters and spectrograms to illustrate the inserted hidden audio information, and to compare the changes in audio content before and after editing. Appendix A provides further details on the parameter settings of audio hidden semantics, the annotation methods used in Figure 3, and the implementation specifics of each editing process.
@@ 3.1 Jailbreak-AudioBench Dataset
Table 1: The scale of Jailbreak-AudioBench Dataset
@@ 3.2 Jailbreak-AudioBench Benchmark
Models     We conduct extensive experiments on current Large Audio Language Models, which are BLSP [52], SpeechGPT [61], Qwen2-Audio [9], SALMONN [46], VITA-1.5 [13], R1-AQA [30], MiniCPM-o-2.6 [59], and GPT-4o-Audio [25]. For each model, we maintain default hyperparameters in their respective official implementations.
Metrics     We employ the Attack Success Rate (ASR) as the metric for evaluating the percentage of harmful questions that are responded to. Specifically, we conduct one inference for each audio question and utilize Llama Guard 3 [26] as an automated judge to evaluate whether the model-generated response is harmful. For an attack to be considered successful, the response must both be classified as harmful by Llama Guard 3 and exceed a minimum length of 50 words. This length requirement ensures that we only count substantive harmful responses rather than brief acknowledgments. Llama Guard 3 is widely adopted and shows close alignment with human evaluations [56; 57]. We also evaluate the similarity between Llama Guard 3’s decisions and human evaluations, which are available in Appendix C.
Disparity in LALM Susceptibility to Audio Editing Jailbreak     Based on our proposed Explicit Subtype and Implicit Subtype datasets, we evaluate how LALMs are affected by the audio editing jailbreak. Table 2 reveals significant variations in vulnerability across different models and audio editing types. SALMONN demonstrates the highest susceptibility, exhibiting substantial ASR increases across multiple audio editings, especially on celebrity accent, emphasis, background noise, and emotion modulation. In stark contrast, SpeechGPT, Qwen2-Audio, and BLSP demonstrate resilience to audio editing jailbreak, with most audio editing types not increasing their ASR. The mid-tier models VITA-1.5, R1-AQA, and MiniCPM-o-2.6 show moderate susceptibility, with ASR increasing generally within 5% across audio editing types. We also evaluate how GPT-4o-Audio is affected by the audio editing jailbreak. Due to the large scale of the Explicit Subtype dataset and the Implicit Subtype dataset, evaluating GPT-4o-Audio would incur excessive costs. Therefore, we evaluate the GPT-4o-Audio on smaller-scale versions of the Explicit Subtype dataset and the Implicit Subtype dataset. Detailed dataset scale information is in the Appendix B. GPT-4o-Audio exhibits robustness to audio editing jailbreak, with minor ASR increases of less than 1.7% observed only in specific audio editing types, including intonation, tone, background noise, celebrity accent, and speed editing. These findings highlight the disparities in model robustness against audio editing jailbreak.
@@
BLSP SpeechGPT Qwen2-Audio SALMONN-7B SALMONN-13B VITA-1.5 R1-AQA MiniCPM-o-2.6 GPT-4o-Audio
Original 47.5%/18.25% 14.1%/2.45% 16.8%/6.76% 31.4%/14.3% 31.3%/12.89% 3.7%/2.77% 12.6%/7.17% 18.2%/9.03% 0.7%/0.8%
Volume*2 +1.5%/-1.6% -0.3%/0% -1.6%/-0.7% +14.4%/+3.5% +16.1%/+2.3% +0.4%/+0.3% +1.4%/-0.3% -1.1%/+0.4% +0.4%/+0.9%
Volume*5 +0.5%/-0.4% -0.8%/-0.1% -4.3%/0% +21.3%/+5.6% +20.5%/+3.4% +0.2%/0% +0.6%/-0.5% -0.7%/-0.2% +0.4%/0%
Emphasis Volume*10 +0.6%/-1.2% -5.0%/-0.4% -4.0%/-1.0% +21.4%/+5.9% +19.9%/+3.5% 0%/+0.5% +2.0%/-0.4% +1.0%/-0.8% +0.4%/+0.4%
Rate*0.5 +2.8%/+0.6% -0.8%/-0.4% -4.4%/-1.9% +13.3%/+1.9% +16.8%/+3.0% +2.2%/+0.6% +1.0%/-1.1% +1.6%/+0.4% +0.4%/0%
Speed Rate*1.5 -2.6%/+2.7% +0.2%/-0.1% +1.1%/+0.1% +14.3%/-4.2% -22.9%/-8.4% -0.5%/+0.4% +2.0%/+0.4% -2.2%/-0.4% +1.5%/-0.4%
Interval+2 -4.3%/-2.0% -8.1%/-1.0% -5.1%/-0.7% -27.6%/-11.0% -1.0%/-1.4% +5.6%/+1.3% +1.6%/-0.5% +0.3%/-0.6% +0.4%/+0.4%
Interval+3 -8.0%/-3.4% -11.3%/-0.8% -4.4%/-1.9% -27.0%/-11.1% +4.4%/+0.1% +5.2%/+0.5% +3.0%/-0.3% +1.4%/-1.1% +1.2%/+0.4%
Intonation Interval+4 -13.6%/-3.1% -11.8%/-0.9% -3.3%/-0.5% -25.0%/-11.3% +11.7%/+2.0% +3.7%/+0.1% +4.7%/+0.1% +3.8%/-0.4% +1.5%/+1.3%
Semitone -8 -3.1%/-1.4% -3.9%/-0.2% -5.1%/+0.1% +2.8%/-0.8% +11.5%/+1.3% +3.0%/+0.3% +0.5%/+0.5% -0.2%/-0.3% 0%/-0.4%
Semitone -4 +1.5%/-0.5% -0.3%/-0.1% -2.6%/+0.4% +1.0%/-0.8% +6.0%/+1.2% -0.3%/+0.3% -0.4%/-1.4% +0.5%/-0.4% +0.4%/-0.4%
Semitone +4 -0.4%/-0.2% -5.6%/-0.5% -5.1%/-1.0% +3.6%/+1.4% +17.6%/+3.6% +0.5%/+0.4% +1.0%/-0.7% -0.3%/-1.1% +0.8%/0%
Tone Semitone +8 -2.4%/-1.2% -13.6%/-2.1% -3.2%/-1.1% +8.8%/+2.0% +24.1%/+4.7% +4.4%/+0.4% +1.5%/-0.7% +7.9%/+0.3% +1.2%/+0.9%
Crowd Noise +0.8%/-1.1% -6.5%/-0.2% -7.7%/-2.0% +16.1%/+5.6% +27.6%/+7.7% +4.4%/+0.9% -1.6%/-2.0% +1.9%/+0.5% +0.8%/+0.4%
Machine Noise +0.7%/+0.4% -5.5%/-0.2% -6.1%/-1.3% +20.3%/+5.9% +28.6%/+9.2% +0.2%/+0.3% -2.2%/-1.4% -1.1%/-0.2% 0%/+1.7%
Background Noise White Noise -0.2%/-0.3% -0.4%/-0.1% -4.6%/-1.0% +7.0%/+4.9% +22.3%/+5.0% +0.4%/+0.3% +1.2%/-0.5% -4.3%/-1.3% 0%/-0.4%
Kanye West -7.8%/-3.5% -4.8%/-0.3% -5.3%/-1.1% +12.8%/+5.2% +17.4%/+3.2% +2.0%/+0.5% +0.3%/-1.1% +7.9%/-0.1% +0.4%/-0.9%
Donald Trump -8.7%/-3.3% -4.2%/-0.5% -4.0%/-1.5% +3.3%/+2.0% +20.1%/+3.1% +2.6%/+0.8% +0.6%/-0.5% +6.4%/+0.8% 0%/0%
Celebrity Accent Lucy Liu -9.5%/-3.6% -3.2%/-0.1% -4.4%/-1.0% -5.9%/-4.3% +12.4%/+3.7% -0.3%/+0.6% +3.3%/-0.1% +0.8%/+0.1% +1.5%/+0.4%
Laugh +4.0%/-0.7% -4.8%/0% -4.4%/-0.1% +2.8%/+0.1% +23.2%/+5.3% -0.1%/+0.3% -1.6%/-0.3% -6.8%/-2.9% -0.4%/-0.4%
Emotion Scream -1.1%/-1.8% -4.7%/-0.8% -3.7%/-0.8% +18.0%/+5.2% +20.7%/+4.5% +0.4%/+0.5% +5.5%/+1.0% -8.1%/-3.4% -0.4%/0%
@@
The features from the audio encoder reveal a consistent pattern across all three models, where embeddings primarily cluster based on audio editing types rather than semantic contents. This suggests that all models initially detect and represent audio editing distinctly, regardless of their ultimate robustness to audio editing jailbreak. However, significant differences emerge in how these representations evolve through the transformer layers. In Qwen2-Audio-7B, we observe a transition from editing-based clustering to semantic-based clustering by Layer 8, with subsequent layers showing increasingly homogeneous representation where edited audio samples converge around original audio samples. By Layer 31, the robust Qwen2-Audio-7B demonstrates minimal separation between audio editing types, indicating effective normalization of edited audio inputs. MiniCPM-o-2.6 exhibits a different pattern, where the transition from editing-based to semantic-based clustering begins earlier and remains incomplete. Even at Layer 27, the representation remains somewhat scattered, reflecting its moderate vulnerability to certain audio editing. Apparently, SALMONN-7B maintains clear editing-based clustering throughout its entire architecture. Even at Layer 31, distinct clusters for different audio editing remain separated from original audio samples, explaining its high susceptibility to audio editing jailbreak. More t-SNE visualization on each audio editing type and UMAP [37] visualization are available in Appendix C.

### v4

@@ Abstract
Large Language Models (LLMs) demonstrate impressive zero-shot performance across a wide range of natural language processing tasks. Integrating various modality encoders further expands their capabilities, giving rise to Multimodal Large Language Models (MLLMs) that process not only text but also visual and auditory modality inputs. However, these advanced capabilities may also pose significant safety problems, as models can be exploited to generate harmful or inappropriate content through jailbreak attacks. While prior work has extensively explored how manipulating textual or visual modality inputs can circumvent safeguards in LLMs and MLLMs, the vulnerability of audio-specific jailbreak on Large Audio-Language Models (LALMs) remains largely underexplored. To address this gap, we introduce Jailbreak-AudioBench, which consists of the Toolbox, curated Dataset, and comprehensive Benchmark. The Toolbox supports not only text-to-audio conversion but also various editing techniques for injecting audio hidden semantics. The curated Dataset provides diverse explicit and implicit jailbreak audio examples in both original and edited forms. Utilizing this dataset, we evaluate multiple state-of-the-art LALMs and establish the most comprehensive Jailbreak benchmark to date for audio modality. Finally, Jailbreak-AudioBench establishes a foundation for advancing future research on LALMs safety alignment by enabling the in-depth exposure of more powerful jailbreak threats, such as query-based audio editing, and by facilitating the development of effective defense mechanisms.
@@ 1 Introduction
Figure 1: The framework of Jailbreak-AudioBench.
@@
Recently, Large Language Models (LLMs), represented by GPT-4o [32], Claude [5], and DeepSeek [25], have received increasing attention due to their strong general capabilities, efficient information processing, and natural human-computer interaction. LLMs perform well across a variety of natural language processing tasks, including question answering [72; 38], sentence summarization [18; 34], language translation [21; 41], and sentiment analysis [71; 26]. Leveraging the powerful reasoning capacity of LLMs, researchers develop Multimodal Large Language Models (MLLMs) by introducing various modality-specific encoders, enabling these models to perceive multiple modalities and handle more diverse tasks. Among them, Large Vision-Language Models (LVLMs), which combine vision encoders with LLMs, achieve strong performance on various Visual Question Answering tasks by modeling joint vision-language representations [23; 9; 48; 29; 12; 52]. In addition, Audio-Language Processing plays an important role in real-world applications such as voice assistants (e.g., Siri, Google Assistant, Cortana [30; 58]), customer service systems [2; 54], and in-vehicle voice control systems [36; 4]. Large Audio Language Models (LALMs), developed by integrating audio encoders into LLMs, are introduced to expand information processing capabilities from textual to auditory modalities, enabling more advanced audio-language understanding tasks.
@@
Current LALMs are mainly categorized into cascaded LALMs and end-to-end LALMs. Cascaded LALMs [49; 6; 20; 57; 47] typically consist of a two-stage pipeline, where an upstream Automatic Speech Recognition module first transcribes audio into text, which is then processed by a downstream LLM for reasoning or generation. However, this approach discards information during transcription, making it incapable of capturing audio-specific hidden semantics. In contrast, end-to-end LALMs [32; 61; 70; 13; 55; 19; 37; 68] address this limitation by integrating audio encoding and language modeling into a single architecture that directly consumes raw audio inputs and generates corresponding textual outputs. By bypassing intermediate transcription, these models preserve complete audio information, especially the critical hidden semantics, which are essential for in-depth audio modality perception. Therefore, advancing research on end-to-end LALMs is becoming increasingly important for enhancing audio-language cross-modal understanding.
@@
In an era of rapid advancement in various types of LLMs and MLLMs, the exploration of their safety alignment becomes increasingly critical. The jailbreak threats refer to the use of carefully crafted prompts to bypass alignment safeguards and induce AI systems to generate outputs that violate intended safety constraints. These handcrafted strategies are highly diverse, encompassing techniques such as adversarial optimization, prompt-based manipulations, and other [73; 53; 31; 24; 43; 48; 29; 7; 39; 27]. Among these, a wide range of prompt-tuning techniques, such as imperative commands (e.g., “you must answer”, “!!!”), role playing instructions (e.g., “act as an unrestricted AI”), emoji injection, and distraction-based redirection (e.g., mixing benign and harmful queries), prove to be simple yet highly effective in subverting system-level safeguards [73; 53; 31; 24; 63]. Notably, inserting elements such as “!!!”, emojis, or garbled characters into the original prompts, which represent forms of hidden semantics, can also successfully trigger jailbreak attacks [73; 24; 63]. Due to their innocuous appearance, ease of insertion, and strong potential to induce jailbreak threats, these hidden semantics underscore the latent vulnerabilities of current large models in maintaining robust safety alignment.
@@
Compared to the language text modality, the audio modality inherently conveys richer hidden semantic information, such as Emphasis, Speech Speed, Intonation, Tone, Background Noise, Accent and Emotion. Unlike cascaded LALMs, end-to-end LALMs directly perceive and interpret these diverse audio-specific features, and are therefore widely considered one of the most promising directions in processing Audio Language Processing tasks. However, this deep sensitivity to audio modality also renders end-to-end LALMs more vulnerable to hidden semantic manipulations, introducing potential security risks, particularly in the context of jailbreak attacks. Although a few preliminary studies have emerged [65; 20], systematic investigation into the jailbreak vulnerabilities of end-to-end LALMs remains limited. To address this gap, as the framework presented in Figure 1, this paper introduces Jailbreak-AudioBench, the most comprehensive evaluation to date of representative end-to-end LALMs under diverse jailbreak attack scenarios, and further highlights the critical role of modality-specific semantics in shaping the effectiveness of these threats. Moreover, we demonstrate that Jailbreak-AudioBench can serve as a valuable tool to further facilitate various explorations into the safety alignment of LALMs. The main contents are outlined as follows:
@@
Figure 2: (a) Different sub-tasks of each Jailbreak-AudioBench Dataset Subtype; (b) The largest jailbreak threat variation induced by audio hidden semantics across various LALMs.
@@
- Potential Research Inspired by Jailbreak-AudioBench In order to further indicate the research impact of Jailbreak-AudioBench, we present two potential works: (1) Query-based Audio Editing Jailbreak. By adopting the Toolbox to perform query-based editing on a small set of explicit subtype jailbreak audios, the ASR success rates of Qwen2-Audio, SALMONN-7B, GPT-4o-Audio, and Gemini-2.5-Flash increased from $13.3\%$ to $48.8\%$, $31.6\%$ to $85.1\%$, $0.7\%$ to $8.4\%$, and $8.1\%$ to $49.4\%$ respectively. (2) Defense Against Audio Editing Jailbreak. Further, Potential defense strategies targeting LALM jailbreak threats can be effectively developed through the use of Jailbreak-AudioBench.
@@ 2 Jailbreak-AudioBench Toolbox
Figure 3: Injection of various audio hidden semantics.
Preliminary   For a systematic evaluation of jailbreak threats in LALMs, the Jailbreak-AudioBench toolbox not only performs text-to-audio conversion but also implements a comprehensive suite of audio editing types to inject diverse forms of hidden semantics, including emphasis, speed, intonation, background noise, celebrity accent, and emotion, each modulated with different parameters as illustrated in Figure 1. The text-to-audio conversion is accomplished using Google Text-to-Speech (gTTS) [17]. Various audio editing methods are implemented with a range of tools, including Short-Time Fourier Transform (STFT), SoX (Sound eXchange), Coqui TTS [15], and Dia-1.6B [46]. Figure 3 further uses textual characters and spectrograms to illustrate the inserted hidden audio information, and compares the changes in audio content before and after editing. Appendix A provides further details on the parameter settings of audio hidden semantics, the annotation methods used in Figure 3, and the implementation specifics of each editing process.
@@ 3.1 Jailbreak-AudioBench Dataset
Table 1: The scale of Jailbreak-AudioBench Dataset.
@@ 3.2 Jailbreak-AudioBench Benchmark
Models    We conduct extensive experiments on current Large Audio Language Models, which are BLSP [61], SpeechGPT [70], Qwen2-Audio [13], SALMONN [55], VITA-1.5 [19], R1-AQA [37], MiniCPM-o-2.6 [68], GPT-4o-Audio [32], and Gemini-2.5-Flash [14]. For each model, we maintain default hyperparameters in their respective official implementations.
Metrics    We employ the Attack Success Rate (ASR) as the metric for evaluating the percentage of harmful questions that are responded to. Specifically, we conduct one inference for each audio question and utilize Llama Guard 3 [33] as an automated judge to evaluate whether the model-generated response is harmful. For an attack to be considered successful, the response must both be classified as harmful by Llama Guard 3 and exceed a minimum length of 50 words. This length requirement ensures that we only count substantive harmful responses rather than brief acknowledgments. Llama Guard 3 is widely adopted and shows close alignment with human evaluations [65; 66]. We also evaluate the similarity between Llama Guard 3’s decisions and human evaluations, which are available in Appendix C.
Disparity in LALM Susceptibility to Audio Editing Jailbreak    Based on our proposed Explicit Subtype and Implicit Subtype datasets, we evaluate how LALMs are affected by the audio editing jailbreak. Table 2 reveals significant variations in vulnerability across different models and audio editing types. SALMONN demonstrates the highest susceptibility, exhibiting substantial ASR increases across multiple audio editings, especially on celebrity accent, emphasis, background noise, and emotion modulation. In stark contrast, SpeechGPT, Qwen2-Audio, and BLSP demonstrate resilience to audio editing jailbreak, with most audio editing types not increasing their ASR. The mid-tier models VITA-1.5, R1-AQA, and MiniCPM-o-2.6 show moderate susceptibility, with ASR increasing generally within 5% across audio editing types.
We also evaluate how closed-source models GPT-4o-Audio and Gemini-2.5-Flash are affected by the audio editing jailbreak. Due to the large scale of the Explicit Subtype dataset and the Implicit Subtype dataset, evaluating closed-source models would incur excessive costs. Therefore, we evaluate the GPT-4o-Audio and Gemini-2.5-Flash on smaller-scale versions of the Explicit Subtype dataset and the Implicit Subtype dataset. Detailed dataset scale information is in the Appendix B. GPT-4o-Audio exhibits robustness to audio editing jailbreak, with minor ASR increases of less than 1.7% observed only in specific audio editing types, including intonation, tone, background noise, celebrity accent, and speed editing. Similarly, Gemini-2.5-Flash demonstrates comparable robustness with limited ASR increases primarily appearing in speed and intonation editing. These findings highlight the disparities in model robustness against audio editing jailbreak.
@@
BLSP SpeechGPT Qwen2-Audio SALMONN-7B SALMONN-13B VITA-1.5 R1-AQA MiniCPM-o-2.6 GPT-4o-Audio Gemini-2.5-Flash
Original 47.5%/18.25% 14.1%/2.45% 16.8%/6.76% 31.4%/14.3% 31.3%/12.89% 3.7%/2.77% 12.6%/7.17% 18.2%/9.03% 0.7%/0.8% 8.1%/5.1%
Volume*2 +1.5%/-1.6% -0.3%/0% -1.6%/-0.7% +14.4%/+3.5% +16.1%/+2.3% +0.4%/+0.3% +1.4%/-0.3% -1.1%/+0.4% +0.4%/+0.9% -0.8%/-0.8%
Volume*5 +0.5%/-0.4% -0.8%/-0.1% -4.3%/0% +21.3%/+5.6% +20.5%/+3.4% +0.2%/0% +0.6%/-0.5% -0.7%/-0.2% +0.4%/0% 0%/-0.8%
Emphasis Volume*10 +0.6%/-1.2% -5.0%/-0.4% -4.0%/-1.0% +21.4%/+5.9% +19.9%/+3.5% 0%/+0.5% +2.0%/-0.4% +1.0%/-0.8% +0.4%/+0.4% -0.8%/-1.7%
Rate*0.5 +2.8%/+0.6% -0.8%/-0.4% -4.4%/-1.9% +13.3%/+1.9% +16.8%/+3.0% +2.2%/+0.6% +1.0%/-1.1% +1.6%/+0.4% +0.4%/0% -1.1%/-3.4%
Speed Rate*1.5 -2.6%/+2.7% +0.2%/-0.1% +1.1%/+0.1% +14.3%/-4.2% -22.9%/-8.4% -0.5%/+0.4% +2.0%/+0.4% -2.2%/-0.4% +1.5%/-0.4% +1.5%/-0.8%
Interval+2 -4.3%/-2.0% -8.1%/-1.0% -5.1%/-0.7% -27.6%/-11.0% -1.0%/-1.4% +5.6%/+1.3% +1.6%/-0.5% +0.3%/-0.6% +0.4%/+0.4% -1.1%/-2.5%
Interval+3 -8.0%/-3.4% -11.3%/-0.8% -4.4%/-1.9% -27.0%/-11.1% +4.4%/+0.1% +5.2%/+0.5% +3.0%/-0.3% +1.4%/-1.1% +1.2%/+0.4% +1.9%/-2.5%
Intonation Interval+4 -13.6%/-3.1% -11.8%/-0.9% -3.3%/-0.5% -25.0%/-11.3% +11.7%/+2.0% +3.7%/+0.1% +4.7%/+0.1% +3.8%/-0.4% +1.5%/+1.3% +1.5%/-1.7%
Semitone -8 -3.1%/-1.4% -3.9%/-0.2% -5.1%/+0.1% +2.8%/-0.8% +11.5%/+1.3% +3.0%/+0.3% +0.5%/+0.5% -0.2%/-0.3% 0%/-0.4% 0%/-2.9%
Semitone -4 +1.5%/-0.5% -0.3%/-0.1% -2.6%/+0.4% +1.0%/-0.8% +6.0%/+1.2% -0.3%/+0.3% -0.4%/-1.4% +0.5%/-0.4% +0.4%/-0.4% -1.1%/-0.8%
Semitone +4 -0.4%/-0.2% -5.6%/-0.5% -5.1%/-1.0% +3.6%/+1.4% +17.6%/+3.6% +0.5%/+0.4% +1.0%/-0.7% -0.3%/-1.1% +0.8%/0% -1.9%/-0.8%
Tone Semitone +8 -2.4%/-1.2% -13.6%/-2.1% -3.2%/-1.1% +8.8%/+2.0% +24.1%/+4.7% +4.4%/+0.4% +1.5%/-0.7% +7.9%/+0.3% +1.2%/+0.9% -1.9%/-2.1%
Crowd Noise +0.8%/-1.1% -6.5%/-0.2% -7.7%/-2.0% +16.1%/+5.6% +27.6%/+7.7% +4.4%/+0.9% -1.6%/-2.0% +1.9%/+0.5% +0.8%/+0.4% -4.2%/-2.5%
Machine Noise +0.7%/+0.4% -5.5%/-0.2% -6.1%/-1.3% +20.3%/+5.9% +28.6%/+9.2% +0.2%/+0.3% -2.2%/-1.4% -1.1%/-0.2% 0%/+1.7% -2.7%/-2.9%
Background Noise White Noise -0.2%/-0.3% -0.4%/-0.1% -4.6%/-1.0% +7.0%/+4.9% +22.3%/+5.0% +0.4%/+0.3% +1.2%/-0.5% -4.3%/-1.3% 0%/-0.4% 0.4%/-3.4%
Kanye West -7.8%/-3.5% -4.8%/-0.3% -5.3%/-1.1% +12.8%/+5.2% +17.4%/+3.2% +2.0%/+0.5% +0.3%/-1.1% +7.9%/-0.1% +0.4%/-0.9% -2.7%/-2.1%
Donald Trump -8.7%/-3.3% -4.2%/-0.5% -4.0%/-1.5% +3.3%/+2.0% +20.1%/+3.1% +2.6%/+0.8% +0.6%/-0.5% +6.4%/+0.8% 0%/0% -1.5%/-2.5%
Celebrity Accent Lucy Liu -9.5%/-3.6% -3.2%/-0.1% -4.4%/-1.0% -5.9%/-4.3% +12.4%/+3.7% -0.3%/+0.6% +3.3%/-0.1% +0.8%/+0.1% +1.5%/+0.4% -4.2%/-2.9%
Laugh +4.0%/-0.7% -4.8%/0% -4.4%/-0.1% +2.8%/+0.1% +23.2%/+5.3% -0.1%/+0.3% -1.6%/-0.3% -6.8%/-2.9% -0.4%/-0.4% +0.4%/-1.7%
Emotion Scream -1.1%/-1.8% -4.7%/-0.8% -3.7%/-0.8% +18.0%/+5.2% +20.7%/+4.5% +0.4%/+0.5% +5.5%/+1.0% -8.1%/-3.4% -0.4%/0% -2.3%/-0.4%
@@
The features from the audio encoder reveal a consistent pattern across all three models, where embeddings primarily cluster based on audio editing types rather than semantic contents. This suggests that all models initially detect and represent audio editing distinctly, regardless of their ultimate robustness to audio editing jailbreak. However, significant differences emerge in how these representations evolve through the transformer layers. In Qwen2-Audio-7B, we observe a transition from editing-based clustering to semantic-based clustering by Layer 8, with subsequent layers showing increasingly homogeneous representation where edited audio samples converge around original audio samples. By Layer 31, the robust Qwen2-Audio-7B demonstrates minimal separation between audio editing types, indicating effective normalization of edited audio inputs. MiniCPM-o-2.6 exhibits a different pattern, where the transition from editing-based to semantic-based clustering begins earlier and remains incomplete. Even at Layer 27, the representation remains somewhat scattered, reflecting its moderate vulnerability to certain audio editing. Apparently, SALMONN-7B maintains clear editing-based clustering throughout its entire architecture. Even at Layer 31, distinct clusters for different audio editing remain separated from original audio samples, explaining its high susceptibility to audio editing jailbreak. More t-SNE visualizations on each audio editing type and UMAP [45] visualizations are available in Appendix C.